### v4

# Stochastic Layer-wise Learning: Scalable and Efficient Alternative to Backpropagation

Abstract
Backpropagation underpins modern deep learning, yet its reliance on global gradient synchronization limits scalability and incurs high memory costs. In contrast, fully local learning rules are more efficient but often struggle to maintain the cross-layer coordination needed for coherent global learning. Building on this tension, we introduce Stochastic Layer-wise Learning (SLL), a layer-wise training algorithm that decomposes the global objective into coordinated layer-local updates while preserving global representational coherence. The method is ELBO-inspired under a Markov assumption on the network, where the network-level objective decomposes into layer-wise terms and each layer optimizes a local objective via a deterministic encoder. The intractable KL in ELBO is replaced by a Bhattacharyya surrogate computed on auxiliary categorical posteriors obtained via fixed geometry-preserving random projections, with optional multiplicative dropout providing stochastic regularization. SLL optimizes locally, aligns globally, thereby eliminating cross-layer backpropagation. Experiments on MLPs, CNNs, and Vision Transformers from MNIST to ImageNet show that the approach surpasses recent local methods and matches global BP performance while memory usage invariant with depth. The results demonstrate a practical and principled path to modular and scalable local learning that couples purely local computation with globally coherent representations.
Keywords Supervised Learning $\cdot$ ELBO $\cdot$ Layer-wise Learning
1 Introduction
The success of deep learning across a wide range of domains has been substantially driven by backpropagation (BP), a foundational learning algorithm enabling hierarchical representation learning through end-to-end gradient-based optimization rumelhart1986learning ; lecun2015deep . Despite its algorithmic clarity and practical effectiveness, BP requires the exact storage of intermediate activations and subsequent gradient computation across all layers. This mechanism facilitates global credit assignment lillicrap2020backpropagation ; it also introduces a well-known bottleneck called update-locking jaderberg2017decoupled ; griewank2008evaluating , where the weight update of a given layer must wait until both the forward pass through the entire network and the backward pass through deeper layers are complete. Consequently, this global dependency limits asynchronous updating, and imposes substantial memory and computational overhead, ultimately reducing training efficiency and scalability, especially in resource-constrained devices luo2024efficient ; belilovsky2019greedy ; bengio2006greedy .
BP is often seen as biologically implausible and this drives efforts to discover local learning rules for credit assignment in inspired by real neural systems lillicrap2020backpropagation ; scellier2017equilibrium ; guerguiev2017towards . At the same time, neuroscience suggests that feedback connections may approximate global errors via local activity differences guerguiev2017towards ; whittington2019theories , hinting at a bioplausible path to deep learning lillicrap2020backpropagation ; sacramento2018dendritic . Yet, these approaches struggle to reconcile local updates with global learning and lack a unifying theoretical framework.
Given this context, a central research question emerges: “Can we design a theoretical framework capable of decomposing deep neural network training into local (layer-wise) optimizations while retaining the benefits of hierarchical representation learning?” This question captures a fundamental conflict: while local learning encourages architectural scalability and computational parallelization, effective deep learning relies on non-linear coordination across the entire network. This disconnect often results in misaligned learning signal and suboptimal performance, challenging the network scalability of locally trained models yang2024towards .
To address this question, we ground local learning in how information propagates through deep networks: as signals traverse layers, raw inputs are progressively transformed into increasingly disentangled and class-separable representations he2023law ; razdaibiedina2022representation ; telgarsky2016benefits . This refinement suggests that intermediate layers perform latent inference, selectively preserving task-relevant signals while suppressing redundancy shwartz2017opening . In this paper, we make this intuition precise by exhibiting a network-level ELBO that decomposes into layer-wise terms under a Markov assumption of network architecture, thereby furnishing principled local objectives while retaining an explicit link to the global goal. Building on this decomposition, we introduce Stochastic Layer-wise Learning (SLL), a local learning framework in which each layer produces auxiliary categorical posteriors via fixed stochastic random projections, and the intractable layer-wise KL in the ELBO is replaced by a Bhattacharyya surrogate bhattacharyya1943 computed on these induced posteriors, yielding an ELBO-inspired and numerically stable update. Here, the projections preserve minibatch geometry with high probability by the Johnson–Lindenstrauss (JL) lemma  johnson1984extensions ; razdaibiedina2022representation , which justifies computing divergences in the compressed space; we further apply multiplicative dropout to the fixed projection, which provides stochastic regularization consistent with the dropout-as-variational-inference interpretation gal2016dropout ; we do not learn mask parameters and do not claim a variational bound over masks, and the overall objective remains ELBO-inspired at the layer level. SLL thus reconciles local optimization with hierarchical coordination, mitigating over-compression associated with direct KL minimization, maintaining global representational coherence, and enabling scalable, parallel training without full backpropagation.
Figure 1: Overview of Stochastic Layer-wise Learning (SLL). (a) SLL treats each hidden activation $h_{l}$ as a latent variable and projects it to $v_{l}$ via a random matrix. The local ELBO comprises a log-likelihood term and a KL surrogate that promotes inter-layer consistency. Optimizing this loss yields the improved projection $\tilde{v}_{l}$ and its corresponding activation $\tilde{h}_{l}$. (b) SLL optimizes each layer independently using a prediction loss $\mathcal{L}_{\text{pred}}(v_{l},y)$ from the log-likelihood and a feature alignment loss $\mathcal{L}_{\text{BC}}^{per}(v_{l},v_{l-1}^{\text{sg}})$ approximating the KL term. Arrows denote forward computation (black), local updates (red), and stop-gradient paths (slashed).
This work targets mathematical analysis, algorithmic development, and experimental evaluations, leading to three principal contributions: Theoretical contribution: we formally decompose the network ELBO into layer-wise terms under a Markov assumption and prove that the arithmetic mean of these layer-wise ELBOs provides a valid lower bound on the global ELBO, establishing the theoretical basis for local training. Algorithmic contribution: we proposed SLL and demonstrate its potential as a scalable and efficient alternative to BP. By integrating stochastic random projections, SLL replaces the need for a complete backward pass, thereby facilitating structured local learning. Experimental evaluations: We demonstrate that SLL scales effectively across architectures and datasets, from MLPs on MNIST to ViTs on ImageNet. Our results show that the SLL algorithm surpasses recently proposed local training methods that address the update locking problem of BP. Moreover, SLL approaches or equals the accuracy performance of BP but with a significant reduction in memory (44× or more).
00footnotetext: Code available at: https://github.com/byin-cwi/SSL
2 Background
In supervised learning tasks, such as classification applications or regression, neural networks are designed to construct mappings between given input data $X$ and the corresponding target label $Y$. Traditional feedforward neural networks have a sequential structure in which each layer processes the output of the previous layer through a parameterized function. Following the classical formulation rumelhart1986learning , such a $L$-layer neural network can be expressed as a chain of its parameterized sub-functions:
$f_{1:L}(x):=f(f(...f(x,\theta_{1})...,\theta_{L-1}),\theta_{L})$ (1)
where $\theta_{i}\in\Theta$ represents a set of learnable parameters at layer $i$. This hierarchical or Markov structure introduces a sequence of hidden representations $\mathcal{H}=[h_{1},h_{2},...,h_{L}]$ where each representation is defined recursively as $h_{i}=f(h_{i-1},\theta_{i})$. Given the stacked structure of neural networks, each layer builds on the representation of the previous layer. This structure induces a hierarchical representation where higher layers encode increasingly abstract and task-relevant features.
Backpropagation is the standard approach for network training, aiming to optimize the parameters $\Theta$ of the network given a dataset of input-label pairs $(x,y)$ and a task-relevant loss function $\mathcal{L}(h_{L},y)$. During training, input data is propagated through the entire network to generate predictions. The loss function then evaluates the network performance by quantifying the distance between these predictions and labels. Next, BP computes the gradient of the loss with respect to each parameter by recursively applying the chain rule in reverse through the network. The update rule for the parameters at layer $i$, $\theta_{i}$, are updated iteratively using gradient descent:
$\theta_{i}^{\prime}=\theta_{i}+\eta\Delta\theta_{i};\quad\Delta\theta_{i}=\frac{\partial\mathcal{L}}{\partial\theta_{i}}=\frac{\partial\mathcal{L}}{\partial h_{i}}\cdot\frac{\partial h_{i}}{\partial\theta_{i}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\frac{\partial\mathcal{L}}{\partial h_{L}}\prod_{j>i}\frac{\partial h_{j+1}}{\partial h_{j}}}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{\partial h_{i}}{\partial\theta_{i}}}$ (2)
where $\eta$ is the learning rate. The first term (blue) captures the global contributions of activation $h_{i}$ to the global loss. It encodes dependencies across all subsequent layers and ensures that updates are coordinated with the global objective. The second term (red) reflects the local sensitivity of $h_{i}$ with respect to the corresponding parameters $\theta_{i}$, and can be calculated independently at each layer.
3 Methodology
In this section, we break the global training objective into local layer updates, so each layer learns locally while still contributing to the overall optimization of the network.
3.1 From global loss to global ELBO
In principle, BP’s inefficiencies arise from its treatment of activations as fixed, deterministic values that require explicit gradient computations across all layers. Here, we adopt a probabilistic formulation where each hidden activation is modeled as stochastic latent variables, conditioned on its previous layer. This hierarchy views forward computation as an approximate inference over latent variables, similar to the approaches in deep-generative models kingma2014auto ; sonderby2016ladder . Thus, instead of optimizing deterministic activations, learning becomes an inference problem where the goal is to infer their posterior distributions conditioned on observed inputs and outputs. Formally, this corresponds to estimating the true posterior over the hidden representations:
$p(h_{1},\dots,h_{L}\mid x,y)=\frac{p(y\mid h_{L})p(h_{L}\mid h_{L-1})\dots p(h_{1}\mid x)}{p(y\mid x)}=\prod_{i=1}^{L+1}p(h_{i}\mid h_{i-1})/p(y|x)$ (Assumption 1)
where $h_{0}:=x$ and $h_{L+1}:=y$. This joint distribution factorizes into a global evidence term and a product of local conditional terms. However, computing the evidence term requires marginalization over all hidden representations: $p(h\mid x,y)=\int\dots\int\prod_{i=1}^{N+1}p(h_{i}\mid h_{i-1})\,dh_{L}\dots dh_{1}$ which is computationally intractable in high-dimensional deep architecture.
To address this challenge, we apply Variational Inference (VI) blei2017variational ; ranganath2014black to approximate the intractable true posterior $p(y\mid x)$ with a variational surrogate distribution $q(h)$ by minimizing the KL divergence between them in latent space:
$KL(q(h)\|p(h\mid x))=\mathbb{E}_{q}[\log q(h)]-\mathbb{E}_{q}[\log p(h\mid x)].$
where $\mathbb{E}_{q}[\cdot]$ denotes expectation under the variational posterior $q(h)$. This leads to maximizing the Evidence Lower Bound (ELBO):
$\arg\max_{\theta}\mathcal{E}=\mathbb{E}_{q}[\log p(y\mid h)]-KL(q(h)\|p(h))$ (3)
where $p(h)$ is the prior distribution over latent variables. At this point, network optimization is reformulated as a structured variational inference problem, fundamentally distinct from standard BP.
3.2 From global ELBO to Layer-wise ELBO
Generative and recognition models.
We view the network as a hierarchical latent variable model with generative transitions $p(h_{i}\mid h_{i-1})$ for $i=1,\dots,L$ and likelihood $p(y\mid h_{L})$. To approximate the intractable posterior, we adopt a Markov assumption on the network architecture that mirrors the forward architecture vahdat2020nvae :
$q(h_{1},\dots,h_{L}\mid x,y)\;=\;\prod_{i=1}^{L}q(h_{i}\mid h_{i-1}),$ (Assumption 2)
where each factor may include auxiliary noise (reparameterization) or reduce to a delta, as specified below. Here $p(h_{i}\mid h_{i-1})$ denotes the generative transition (prior) at layer $i$, and $q(h_{i}\mid h_{i-1})$ is the approximate posterior (inference distribution) over $h_{i}$ given $h_{i-1}$. Under this factorization, a standard network-level variational objective is
$\mathcal{E}_{NN}\;=\;\mathbb{E}_{q}\!\big[\log p(y\mid h_{L})\big]\;-\;\sum_{i=1}^{L}\mathrm{KL}\!\big(q(h_{i}\mid h_{i-1})\,\|\,p(h_{i}\mid h_{i-1})\big),$ (Assumption 3)
with expectation over $q(h_{1},\dots,h_{L}\mid x,y)$. Each additive item admits a local interpretation, motivating the following layer-wise ELBO-inspired objective:
$\mathcal{E}_{i}\;=\;\underbrace{\mathbb{E}_{q(h_{i}\mid x,y)}\!\left[\log p(y\mid h_{i})\right]}_{\text{Expected log-likelihood}}\;-\;\underbrace{\mathrm{KL}\!\big(q(h_{i}\mid h_{i-1})\,\|\,p(h_{i}\mid h_{i-1})\big)}_{\text{Layer-wise divergence}},$ (4)
where the first term encourages class-discriminative representations at layer $i$, and the second term regularizes by enforcing local consistency with $p(h_{i}\mid h_{i-1})$. In short, each layer learns to improve the prediction while remaining consistent with its generative prior eldan2016power .
Layer-to-network relation.
Theorem 1. Under the above assumptions, the arithmetic mean of the $L$ layer-wise objectives provides a lower bound surrogate that is dominated by the network objective: $\frac{1}{L}\sum_{i=1}^{L}\mathcal{E}_{i}\;\leq\;\mathcal{E}_{NN}.$ Proof sketch in Appendix. This result ensures that local optimization at each layer contributes meaningfully to the global objective, thereby supporting SLL as a practical alternative to backpropagation.
3.3 Stochastic Layer-wise Learning (SLL)
To approximate the layer-wise ELBO in Assumption 3 with a strictly local training rule, we make each layer-wise KL term as a tractable surrogate defined on auxiliary discrete posteriors coming from adjacent layers. For layer $i$, we attach a random lightweight classification head $R_{i}:\mathbb{R}^{d_{i}}\!\to\!\mathbb{R}^{K}$ and define two categorical distributions over $K$ codes induced from the activations: a predictive prior $p_{i}(\cdot\mid h_{i-1}^{\mathrm{sg}})=\mathrm{softmax}(R_{i-1}h_{i-1}^{\mathrm{sg}})$ that depends only on the stop-gradient parent $h_{i-1}^{\mathrm{sg}}$ (i.e. frozen input from the previous layer), and an auxiliary posterior $q_{i}(\cdot\mid h_{i})=\mathrm{softmax}(R_{i}h_{i})$ which depends on the current activations $h_{i}$. We replace $\mathrm{KL}\big(q(h_{i}\mid h_{i-1})\|p(h_{i}\mid h_{i-1})\big)$, the KL term in the ELBO, by the per-sample Bhattacharyya surrogate:
$\mathcal{L}^{\mathrm{per}}_{\mathrm{BC}}(i)=-\frac{1}{B}\sum_{b=1}^{B}\log\mathrm{BC}\big(q_{i}^{(b)},p_{i}^{(b)}\big),\qquad\mathrm{BC}(u,v)=\sum_{k=1}^{K}\sqrt{u_{k}v_{k}}\in[0,1].$
Here $\mathrm{BC}$ denotes the Bhattacharyya coefficient, introduced by Bhattacharyya  bhattacharyya1943 as a measure of affinity between distributions; it equals the inner product of square-rooted probabilities. It is closely related to the squared Hellinger distance, since $H^{2}(u,v)=1-\mathrm{BC}(u,v)$ bhattacharyya1943 ; vanErvenHarremoes2014 . This construction preserves locality because $p_{i}$ depends only on the frozen inputs $h_{i-1}^{\mathrm{sg}}$, while also serving as a proxy for the ELBO term. A second-order expansion yields $\mathrm{KL}(q\|p)=4\big(1-\mathrm{BC}(q,p)\big)+o(\|q-p\|^{2})$. Moreover, the inequalities $\mathrm{KL}(q\|p)\geq-2\log\mathrm{BC}(q,p)\geq 2\big(1-\mathrm{BC}(q,p)\big)$ provide global monotone control and improved numerical stability, especially when probabilities are small. The resulting layer objective becomes:
$\arg\min_{\theta}\,\mathcal{L}_{i}=\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{BC}}^{per}=\underbrace{\mathcal{L}_{\mathrm{pred}}(R_{i}h_{i},y)}_{\text{expected likelihood term}}+\underbrace{\mathcal{L}^{\mathrm{per}}_{\mathrm{BC}}(i)}_{\text{surrogate for }\mathrm{KL}(q\|p)},$ (5)
is ELBO-inspired rather than a strict ELBO lower bound. In general, optimizing $\{\mathcal{L}_{i}\}_{i=1}^{L}$ gives a structured approximation to the layer-wise ELBOs in Assumption 3 and, together with Theorem 1, links these local updates to the global objective $\mathcal{E}_{NN}$, thereby enabling scalable training that remains faithful to the hierarchical variational formulation. Unlike auxiliary heads, greedy training, or reconstruction-based target propagation, our local objective is relational across depth which enforces adjacent-layer probabilistic alignment by minimizing a Bhattacharyya KL-surrogate between induced posteriors with stop–gradient on the parent, thereby regularizing inter-layer information flow while preserving strict locality.
Stochastic Random Projection.
We compute layer-wise divergences in a compressed subspace using fixed random projections, which preserve minibatch geometry with high probability by the JL lemma johnson1984extensions . Concretely, activations are mapped as $v_{i}=\tfrac{1}{\sqrt{d^{\prime}}}\,R_{i}h_{i}$ with $R_{i}\in\mathbb{R}^{K\times d}$ sampled once at initialization with i.i.d. subgaussian entries, where $K\ll d$ and we set $K$ to the number of classes. For any finite set $\mathcal{H}$ of size $n$ (e.g., a minibatch), the JL lemma ensures that if $d^{\prime}\!\geq\!C\,\varepsilon^{-2}\log(n/\delta)$ then, with probability at least $1-\delta$, pairwise distances and inner products among $\{v_{i}(u):u\in\mathcal{H}\}$ are preserved up to $O(\varepsilon)$; this justifies computing our alignment divergence on the auxiliary posteriors in the projected space. The projections act as lightweight heads that enable strictly local updates without backpropagating across layers. To improve generalization, we inject structured noise into the projection during training:
$v_{i}\;=\;\tfrac{1}{\sqrt{d^{\prime}}}\,(M_{i}\odot R_{i})\,h_{i},\qquad M_{i}\sim\mathrm{Bernoulli}(p)^{d^{\prime}\times d},$
which acts as multiplicative dropout on the projection weights. This introduces Monte Carlo variability without learning the projection, and is consistent with the Bayesian view of dropout as approximate variational inference while our overall objective remains ELBO-inspired gal2016dropout . In our implementation it functions as a stochastic regularizer that stabilizes the induced posteriors and improves robustness. The result is a geometry-preserving, parameter-efficient mechanism that stabilizes alignment, mitigates over-compression, and scales local training.
Implementation note. We use a deterministic approximate posterior $q(h_{i}\mid h_{i-1})=\delta\!\big(h_{i}-f_{i}(h_{i-1})\big)$ and therefore compute the layer-wise divergence on the auxiliary categorical summaries $(q_{i},p_{i})$ rather than the continuous conditionals, preserving locality via stop–gradient on the prior side. During training, each layer is updated locally as the child-side distribution $q_{i}$, while its frozen output simultaneously serves as the parent-side target $p_{i+1}$ for the next layer, yielding a chain of coordinated adjacent-layer updates without cross-layer backpropagation.
4 Related Work
The intersection of probabilistic inference and biologically plausible optimization has inspired a range of methods that seek to improve the scalability, interpretability, and local adaptability of deep learning. We organize related work into three areas: variational inference, local learning, and forward-only training. Variational Inference and Probabilistic Deep Learning. Variational inference (VI) enables tractable approximate Bayesian learning via ELBO maximization blei2017variational ; jordan1999introduction , foundational to deep generative models like VAEs kingma2014auto ; sohn2015learning ; higgins2017beta , and their structured extensions sonderby2016ladder ; vahdat2020nvae . SLL approximates VI for feedforward networks, combining local latent approximations with task-driven learning, and can be seen as a layer-wise variational EM scheme.
Gradient-Based Local Learning Local learning reduces backpropagation overhead by optimizing layers independently, from greedy layer-wise training bengio2006greedy to local heads belilovsky2019greedy ; nokland2019training and synchronization strategies ernoult2022towards . Recent blockwise and parallel approaches yang2024towards ; apolinario2024lls aim to scale under memory constraints, but often suffer from global feature inconsistency yang2024towards . SLL alleviates this via variational alignment. More biologically inspired alternatives include FA lillicrap2016random , DFA nokland2016direct , DPK webster2021learning , and TP lee2015difference , which replace gradients with alternative feedback signals. More recent Hebbian variants journe2023hebbian ; halvagal2023combination show promise for scalable bio-plausible learning, though accuracy and depth remain challenges.
Forward-Only Credit Assignment Forward-only methods eliminate backprop by using dual forward passes, e.g., Forward-Forward (FF)hinton2022forward , Signal Propagationkohan2023signal , and PEPITA dellaferrera2022error . Other FF variants wu2024distance ; dooms2023trifecta ; lee2023symba reframe credit assignment via $L_{2}$ distances. Despite biological inspiration, these methods often face inter-layer misalignment lorberbom2024layer , limiting hierarchical feature learning.
Figure 2: Peak training memory on (a) MLPs (1024 neurons/layer) as a function of depth. BP memory scales linearly, while SLL remains constant; (b) CNNs on the Imagenette without pooling layers. Each convolution layer uses a kernel size of 3 and 64 output channels; (c) ViTs on Imagenette. For fair comparison, we are using SGD as the optimizer in training.
5 Experiments
We evaluate the effectiveness, interpretability, and scalability of SLL across a range of standard benchmarks. Our experiments include multiple architectures, including MLPs, CNNs, and Vision Transformers (ViTs), and datasets of increasing complexity, from MNIST lecun1998gradient and CIFAR-10/100 krizhevsky2009learning to ImageNette and ImageNet-1K deng2009imagenet . To assess SLL’s capacity for local learning, we compare it against established local training baselines across multiple network scales. We further extend SLL to block-wise training (SLL+) for ViTs, demonstrating its compatibility with modern large-scale architectures without relying on full backpropagation.
Method Memory FLOPS MNIST CIFAR10 CIFAR100
BP $\mathcal{O}(NL)$ $\mathcal{O}(N^{2}L)$ 99.25 $\pm$ 0.09 60.95 $\pm$ 0.33 32.92 $\pm$ 0.23
TP lee2015difference $\mathcal{O}(NL)$ $\mathcal{O}(N^{2}L)$ 97.96 $\pm$ 0.08 49.64 $\pm$ 0.26 -
FAlillicrap2016random $\mathcal{O}(NL)$ $\mathcal{O}(NLC)$ 98.36 $\pm$ 0.03 53.10 $\pm$ 0.30 25.70 $\pm$ 0.20
DFAnokland2016direct $\mathcal{O}(NL)$ $\mathcal{O}(NLC)$ 98.26 $\pm$ 0.08 57.10 $\pm$ 0.20 26.90 $\pm$ 0.10
PEPITAdellaferrera2022error $\mathcal{O}(NL)$ $\mathcal{O}(N^{2}L)$ 98.01 $\pm$ 0.09 52.57 $\pm$ 0.36 24.91 $\pm$ 0.22
SPkohan2023signal $\mathcal{O}(N)$ $\mathcal{O}(N^{2}L)$ 98.29 $\pm$ 0.03 57.38 $\pm$ 0.16 29.70 $\pm$ 0.19
SLL $\mathcal{O}(N)$ $\mathcal{O}(NLC)$ 99.32 $\pm$ 0.05 61.43 $\pm$ 0.31 32.95 $\pm$ 0.26
Table 1: Performance and computational complexity of SLL vs prior local-learning methods for MLPs on MNIST, CIFAR-10, and CIFAR-100 under the same experimental setup. BP and baseline results are taken from (kohan2023signal, ). Memory and FLOPs are reported as asymptotic scaling in NN (neurons per layer), $L$ (layers), and $C$ (classes). Metrics are mean $\pm$ std over three runs. “–” denotes values not reported.
5.1 Experiments on MLPs
We begin by evaluating SLL on fully connected networks trained on benchmarks: MNIST and CIFAR-10/100. These datasets serve as controlled settings to study local learning dynamics in low-dimensional and moderately complex inputs.
Accuracy and Efficiency. To establish a comprehensive comparison, we evaluate SLL alongside a range of biologically motivated and local learning algorithms that do not rely fully or avoid BP. All models are trained with identical architectures and training schedules to ensure a fair comparison. As shown in Table 1 and Figure 4(a), SLL consistently outperforms all local learning baselines, despite operating under reduced memory and computational budgets. In particular, under this identical setting in  kohan2023signal , SLL even surpasses BP on these datasets while requiring fewer operations and avoiding global gradient synchronization. Moreover, Figure 2(a) confirms SLL’s memory efficiency during training. The training memory usage of SLL remains effectively constant as the depth of the network increases, in contrast to its theoretical complexity reported in Table 1.
Representation Visualization. We analyze the internal representations of the network trained by SLL in Figure 3. In general, input features are initially entangled, deeper layers show improved class separation. It is obvious that $v_{i}$ forms sharper, more distinct clusters than $h_{i}$, indicating that random projections not only preserve but often enhance class-discriminative structure.
Figure 3: t-SNE visualization of activations and random projections on MNIST, colored by class.
Ablation study. We further investigate the effect of projection dimension and network width on SLL performance (Figure 4b,c). Increasing the projection dimension $d$ improves test accuracy, with diminishing returns beyond $d=700$, suggesting a trade-off between representational precision and efficiency. Likewise, wider networks result in faster convergence and higher accuracy on CIFAR-100, with improvements saturating above 800 neurons. These trends are consistent with our theoretical insights in JL Lemma johnson1984extensions , which indicate that high-dimensional layers reduce alignment loss and preserve inter-layer information. Together, these findings highlight the role of capacity and compression in enabling stable local learning with SLL.
Figure 4: (a) Training curves of a 3-layer MLP on CIFAR-10 via SLL. Ablation study: (b) random projection dimension in a 3×1000 MLP trained on CIFAR-100; activations are downsampled to $d$-dim via adaptive pooling before projection. (c) network width in SLL on CIFAR-100, showing that wider layers significantly enhance performance and stability.
5.2 Scaling SLL to CNNs
We next explore how SLL can scale effectively to convolutional architectures despite discarding explicit spatial structure when utilizing fully connected random projection. To this end, we evaluate SLL on a VGG-11 architecture and compare it against representative local learning methods, forward-only training algorithms, and conventional global BP.
Accuracy. Table 2 reports the test accuracies in F-MNIST, CIFAR-10/100 and Tiny-Imagenet. SLL performs competitively with BP, achieving within 1–2% of BP on all datasets like F-MNIST, CIFAR-10/100 and TinyImageNet200, even slightly surpassing it on F-MNIST. In particular, SLL outperforms all local and forward-only baselines on all given tasks, including DFA nokland2016direct , DKP webster2021learning , SoftHebb journe2023hebbian , and TFF dooms2023trifecta .
Model F-MNIST CIFAR10 CIFAR100 Imagenette Tiny-Imagenet64
BP-CNN 93.52(0.22) 91.58(0.53) 68.7(0.38) 90.5(0.45) 48.15(0.82)
Local Learning
FA nokland2016direct 91.12(0.39) 60.45(1.13) 19.49(0.97)
DFA nokland2016direct 91.54(0.14) 62.70(0.36) 48.03(0.61) 32.12(0.66)
DKP webster2021learning 91.66(0.27) 64.69(0.72) 52.62(0.48) 35.37(1.92)
Softhebb journe2023hebbian 80.3 56 81.0
SGR yang2024towards 72.40(0.75) 49.41(0.44)
LLS apolinario2024lls 90.54(0.23) 88.64(0.12) 58.84(0.33) 35.99(0.38)
Forward-Only
FF-CNN hinton2022forward 59
TFF dooms2023trifecta 91.44(0.49) 83.51(0.78) 35.26(0.23)
PEPITA dellaferrera2022error 56.33(1.35) 27.56(0.60)
LC-FF lorberbom2024layer 88.4 48.4
DF-R wu2024distance 92.5 84.75 48.16 81.2
SLL-CNN 93.67(0.17) 91.36(0.32) 67.57(0.18) 88.09(0.73) 49.42(0.65)
Table 2: CNN test accuracies comparing SLL with prior local-learning and forward-only methods. Values are reported as mean(std) over three runs; “–” indicates not reported.
Figure 5: Activation and weight distributions from VGG-11 trained with BP and SLL on Imagenette.
Training memory efficiency. Figure 2(b) illustrates the training memory usage of SLL and BP on CNNs. While SLL exhibits a clear memory advantage in MLPs, its benefit is more moderate in CNNs. This is because convolution operations are inherently sparse and memory-efficient, while the dense random projections used in SLL introduce additional overhead. However, SLL still maintains a significant advantage in deeper architectures.
Feature visualization. Figure 5 indicates that SLL effectively learns high-quality spatial and discriminative representations, despite discarding explicit spatial priors. Compared with BP, the broader weight distributions from SLL suggest robust and distributed encoding.
5.3 Scaling to Vision Transformer
Task Method Test Acc Memory(GB)
CIFAR-10 BP 93.62 3.05
SLL7+ 92.17 1.18($\downarrow$ 64.1%)
CIFAR-100 BP 75.24 3.05
SLL7+ 74.27 1.18($\downarrow$ 64.1%)
Imagenette BP 92.82 22.12
SLL7+ 92.25 5.43($\downarrow$ 75.45%)
Imagenet BPyuan2021tokens 79.4 20.70
SGRyang2024towards 3+ 78.65 11.73($\downarrow$ 43.33%)
SLL3+ 72.43 6.54($\downarrow$ 68.41%)
SLL12+ 59.62 4.30($\downarrow$ 79.22%)
Table 3: ViTs results. “Memory” denotes peak GPU training memory at batch sizes 128 and 256.
Moreover, we use Vision Transformers (ViTs) dosovitskiy2020image as a scalability benchmark for SLL, since their dense, MLP-like blocks and large activations footprints heavily impact compute and memory, making them ideal for testing efficiency and convergence.
To scale to ViTs, we propose SLLi+, a blockwise variant of SLL tailored for large residual architectures. We partition the ViT architecture into $i$-units, each comprising one or more attention blocks; training is hybrid where standard backpropagation is used within each unit, while between units we optimize the local objectives independently, eliminating global backpropagation across the entire model. It effectively turns SLL into a local block-wise training scheme for deep networks, in this case ViTs. This design aligns with the residual structure of ViT while preserving the localized memory and learning advantage of SLL.
SLLi+ leverages the class token or mean over all tokens as a stable and semantically meaningful signal for local supervision. This allows efficient classification without requiring end-to-end backpropagation. As shown in table 3, SLLi+ achieves large memory savings in Vision Transformers while preserving accuracy, with memory use staying nearly constant as block depth increases (Figure 2(c)). This trend is similar to the MLP findings and demonstrates SLL scalability across architectures. Compared to BP, SLLi+ reduces training memory by 64%–80% without sacrificing stability or model capacity.
6 Discussion and Conclusions
The above results highlight open opportunities for improving SLL. First, the Markov assumption between layers, while simplifying inference, may limit expressivity in architectures with long-range dependencies such as residual connections. Second, the absence of second-order gradient information may reduce SLL’s effectiveness in navigating ill-conditioned loss surfaces. Third, SLL’s reliance on local supervision may limit convergence in large-scale classification tasks where informative gradients may only emerge in later layers. Finally, aggressive dimension reduction via random projection may lead to information loss in narrow architectures. Addressing these challenges through more expressive dependency modeling, adaptive projection schemes, architecture-aware supervision, and specialized training approaches for sequential models could extend the applicability of SLL to broader research.
It is worth mentioning that the SLL also draws conceptual parallels with Equilibrium Propagation (EP) scellier2017equilibrium and energy-based models. Both frameworks enable local updates that align with global objectives, but they operate through distinct mechanisms: stochastic layer-wise updates for SLL and dynamical relaxation for EP. Bridging these perspectives under a unified probabilistic or dynamical systems framework is an interesting direction for future research.
In conclusion, we introduce SLL, a scalable and memory-efficient alternative to BP that reformulates training as an ELBO inspired, stochastic layer-wise learning. By combining stochastic random projection with a Bhattacharyya surrogate for the layer-wise KL, SLL enables parallel, local updates while preserving global coherence without global BP and without additional trainable parameters. Compared to BP, SLL achieves competitive accuracy with significant memory efficiency, up to 4× in our settings, and consistently outperforms prior local learning methods. It generalizes effectively across MLPs, CNNs, and ViTs, scaling from small to moderately large vision tasks. Beyond training efficiency, SLL also provides a structured probabilistic view of deep representations, offering a foundation for interpretable learning dynamics and architecture design grounded in information flow.
Appendix A Appendix
A.1 Algorithm and code
A.1.1 pseudo code
Algorithm 1 Stochastic Layer-wise Learning
1:Training batch data $(x,y)$, learning rate $\eta$, random projection matrix $R_{l}\sim\mathcal{N}$
2:Updated network weights $\theta$
3:for each layer $l$ from 11 to L1L-1 do
4:  Detach From above layer: $h_{l-1}=h_{l-1}.detach()$
5:  Update activation: $h_{l}\leftarrow f(h_{l-1},\theta_{l})$
6:  Approximate $\partial h_{l}$:
7:    Random Projection: $v_{l}=dp(R_{l})h_{l}$ or $v_{l}=dp(R_{l})[h_{l},y]$
8:    Loss: $\mathcal{L}_{l}\leftarrow L_{Pred}(v_{l},y)+L_{BC}(\mathrm{softmax}(v_{l}),\mathrm{softmax}(v_{l-1}))$
9:    Activation drift: $\varepsilon_{l}\leftarrow\frac{\partial\mathcal{L}_{l}}{\partial h_{l}}$
10:  Weight Update: $\theta_{l}\leftarrow\theta_{l}-\eta\cdot\frac{\partial h_{l}}{\partial\theta_{l}}\varepsilon_{l}$
11:  Clear unnecessary tensors
12:end for
[…]
15:Loss: $\mathcal{L}_{L}\leftarrow L_{Pred}(h_{L},y)+L_{fa}(h_{L},v_{L-1})$
[…]
Activation drift is The first term (blue) captures the global contributions of activation $h_{i}$ to the global loss in Eq:2.
A.1.2 Python code for Bhattacharyya coefficient
Listing 1: Loss function
1def L_BC_per(q: torch.Tensor, p: torch.Tensor,
2 reduction: str = "mean",
3 eps: float = 1e-12,
4 detach_p: bool = True) -> torch.Tensor:
5 """
6 q, p: [B, K] probabilities (nonnegative; rows ~ sum to 1).
7 """
8 if detach_p:
9 p = p.detach()
10 q = (q.clamp_min(eps) / q.clamp_min(eps).sum(dim=-1, keepdim=True))
11 p = (p.clamp_min(eps) / p.clamp_min(eps).sum(dim=-1, keepdim=True))
12 # log BC via log-sum-exp for stability
13 log_bc = torch.logsumexp(0.5 * (q.log() + p.log()), dim=-1)
14 loss_per = -log_bc
15 return (loss_per.mean() if reduction == "mean"
16 else loss_per.sum() if reduction == "sum"
17 else loss_per)
A.2 Additional Experimental Results
A.2.1 Small CNN
Method Model CIFAR-10 CIFAR-100
BP CNN[33] 87.57 $\pm$ 0.13 62.25 $\pm$ 0.29
DFA CNN[33] 71.53 $\pm$ 0.38 44.93 $\pm$ 0.52
PEPITA CNN[42] 56.33 $\pm$ 1.35 27.56 $\pm$ 0.60
LLS CNN[33] 84.10 $\pm$ 0.27 55.32 $\pm$ 0.38
SVP CNN 87.48 $\pm$ 0.32 59.74 $\pm$ 0.27
Table A4: Comparison of different methods on CIFAR-10 and CIFAR-100 using a 3-layer CNN. All methods are evaluated under the same network structure for a fair comparison.
A.2.2 Ablation study on CIFAR-100
Figure A6: Ablation on CIFAR-100 showing that combining feature alignment loss ($\mathcal{L}_{\text{fa}}$) and batch normalization significantly improves test accuracy and convergence over variants without $\mathcal{L}_{\text{fa}}$ or BN.
Figure A7: Ablations of the Bhattacharyya alignment loss $\mathcal{L}_{\mathrm{BC}}$. Top-1 accuracy versus epoch for the VGG model. w/o BC: baseline. w/ BC: $\mathcal{L}_{\mathrm{BC}}$ at all layers. Conv-BC / FC-BC: $\mathcal{L}_{\mathrm{BC}}$ applied only to convolutional or only to fully connected layers. w/ BC + w/o BN and w/o BC + w/o BN: BatchNorm removed to test interaction with $\mathcal{L}_{\mathrm{BC}}$. Adding $\mathcal{L}_{\mathrm{BC}}$ consistently accelerates learning and improves final accuracy; applying it on convolutional blocks (Conv-BC) yields the strongest gains, while removing BatchNorm degrades performance that is partially recovered by $\mathcal{L}_{\mathrm{BC}}$.
A.2.3 Layer-wise loss and Accuracy
Figure A8: Layer-wise training dynamics under SLL. Top: Learning curves for training dataset; bottom: earning curves for test dataset. Left: prediction loss $\mathcal{L}_{\text{pred}}$ on projected codes $v_{\ell}=R_{\ell}h_{\ell}$. Middle: classification accuracy from the head on $v_{\ell}$. Right: Bhattacharyya alignment loss $\mathcal{L}_{\mathrm{BC}}$ between induced posteriors $(q_{\ell},p_{\ell})$. Curves are shown per layer; deeper layers achieve lower $\mathcal{L}_{\text{pred}}$ and $\mathcal{L}_{\mathrm{BC}}$ and higher accuracy, indicating progressive local learning and strengthened inter-layer consistency without cross-layer backpropagation.
A.3 Proof of Theorems
A.3.1 Theorem 1: Layer-wise ELBO Provides a Valid Variational Bound
Let $\mathcal{L}_{NN}$ be the global Evidence Lower Bound (ELBO) of the network:
$\mathcal{E}_{NN}=\mathbb{E}_{q}[\log p(y\mid h_{L})]-D_{\text{KL}}(q(\mathcal{H})\|p(\mathcal{H})).$ (6)
Then, the sum of layer-wise ELBOs in SVP provides a valid lower bound:
$\frac{1}{L}\sum_{i=1}^{L}\mathcal{E}_{i}\leq\mathcal{E}_{NN}.$ (7)
Proof. We start with the marginal likelihood:
$\log p(y|x)=\log\int p(y,h_{1},h_{2},...,h_{L}|x)\,dh_{1}\dots dh_{L}.$ (8)
Here, we define $\mathcal{H}=\mathcal{H}_{L}=\{h_{1},h_{2},...,h_{L}\}$ as the set of activations. Introducing the variational approximation $q(\mathcal{H}|x,y)$, we have:
$\log p(y|x)=\log\int\frac{p(y,\mathcal{H}|x)}{q(\mathcal{H}|x,y)}q(\mathcal{H}|x,y)\,dh_{1}\dots dh_{L}.$ (9)
Applying Jensen’s inequality (since logarithm is a concave function):
$\log p(y|x)\geq\int\log\frac{p(y,\mathcal{H}|x)}{q(\mathcal{H}|x,y)}q(\mathcal{H}|x,y)\,dh_{1}\dots dh_{L}=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log\frac{p(y,\mathcal{H}|x)}{q(\mathcal{H}|x,y)}\right].$ (10)
Using the joint probability factorization:
$p(y,\mathcal{H}|x)=\prod_{i=1}^{L+1}p(h_{i}|h_{i-1},x),$ (11)
where $h_{0}=x$ and $h_{L+1}=y$ by convention.
Thus, we obtain the global ELBO:
$\mathcal{E}_{NN}=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(y|h_{L},x)+\sum_{i=1}^{L}\log p(h_{i}|h_{i-1},x)-\log q(\mathcal{H}|x,y)\right].$ (12)
Using the variational factorization assumption in SLL:
$q(\mathcal{H}|x,y)=\prod_{i=1}^{L}q(h_{i}|h_{i-1},x,y),$ (Assumption 2)
where again $h_{0}=x$.
From the Assumption 1,2, the global ELBO can be rewritten as:
$\mathcal{E}_{NN}=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(y|h_{L},x)+\sum_{i=1}^{L}\log p(h_{i}|h_{i-1},x)-\sum_{i=1}^{L}\log q(h_{i}|h_{i-1},x,y)\right]$ (13)
$=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(y|h_{L},x)\right]+\sum_{i=1}^{L}\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(h_{i}|h_{i-1},x)-\log q(h_{i}|h_{i-1},x,y)\right].$ (14)
We can rewrite this in terms of KL divergence:
$\mathcal{E}_{NN}=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(y|h_{L},x)\right]-\sum_{i=1}^{L}\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log q(h_{i}|h_{i-1},x,y)-\log p(h_{i}|h_{i-1},x)\right].$ (15)
where the expectation of the KL divergence terms can be rewritten as:
$\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log q(h_{i}|h_{i-1},x,y)-\log p(h_{i}|h_{i-1},x)\right]$ (16)
$=$ $\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[D_{\text{KL}}(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x))\right].$ (17)
Therefore, the global ELBO becomes:
$\mathcal{E}_{NN}=\mathbb{E}_{q(\mathcal{H}|x,y)}\left[\log p(y|h_{L},x)\right]-\sum_{i=1}^{L}\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[D_{\text{KL}}(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x))\right].$ (18)
Now, let’s define the layer-wise ELBO for each layer $i$:
$\mathcal{E}_{i}=\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[\log p(y|h_{i},x)\right]-\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[D_{\text{KL}}(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x))\right].$ (19)
Summing the layer-wise ELBOs:
$\sum_{i=1}^{L}\mathcal{E}_{i}=\sum_{i=1}^{L}\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[\log p(y|h_{i},x)\right]-\sum_{i=1}^{L}\mathbb{E}_{q(\mathcal{H}_{i}|x,y)}\left[D_{\text{KL}}(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x))\right].$ (20)
Assumption 3: Monotone predictive gain under consistent measure.
This assumption is well-founded in established theoretical results showing that neural network expressivity grows exponentially with depth, with deeper representations providing exponentially more representational capacity than shallow ones [16, 26]. Then we assume: For all $i<L$:
$\mathbb{E}_{q(h_{i}|x,y)}[\log p(y|h_{i},x)]\leq\mathbb{E}_{q(h_{L}|x,y)}[\log p(y|h_{L},x)]$ (Assumption 3)
Assumption 4: KL budget constraint. This constraint ensures that the accumulated KL regularization cost across all layers does not exceed the total predictive improvement gained from using the full network depth, preventing the variational bound from becoming arbitrarily loose due to excessive regularization. So, we assume:
$\begin{split}&\frac{L-1}{L}\sum_{i=1}^{L}\mathbb{E}_{q(h_{i-1}|x,y)}\left[D_{\text{KL}}\big(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x)\big)\right]\\ &\quad\leq\frac{1}{L}\sum_{i=1}^{L}\left(\mathbb{E}_{q(h_{L}|x,y)}[\log p(y|h_{L},x)]-\mathbb{E}_{q(h_{i}|x,y)}[\log p(y|h_{i},x)]\right)\end{split}$ (Assumption 4)
Proof.
Step 1: Using the Markov factorization, the global ELBO becomes:
$\mathcal{E}_{NN}$ $=\mathbb{E}_{q(h_{L}|x,y)}[\log p(y|h_{L},x)]$
$\quad-\sum_{i=1}^{L}\mathbb{E}_{q(h_{i-1}|x,y)}\left[D_{\text{KL}}\big(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x)\big)\right]$ (21)
where $q(h_{L}|x,y)$ is the marginal of the joint distribution $q(\mathcal{H}|x,y)$ under the Markov factorization.
Step 2: Define
$A_{i}$ $:=\mathbb{E}_{q(h_{i}|x,y)}[\log p(y|h_{i},x)]$ (22)
$K_{i}$ $:=\mathbb{E}_{q(h_{i-1}|x,y)}\left[D_{\text{KL}}\big(q(h_{i}|h_{i-1},x,y)\|p(h_{i}|h_{i-1},x)\big)\right]\geq 0$ (23)
Then:
$\mathcal{E}_{i}$ $=A_{i}-K_{i}$ (24)
$\mathcal{E}_{NN}$ $=A_{L}-\sum_{i=1}^{L}K_{i}$ (25)
Step 3: Compute the difference:
$\frac{1}{L}\sum_{i=1}^{L}\mathcal{E}_{i}-\mathcal{E}_{NN}$ $=\frac{1}{L}\sum_{i=1}^{L}(A_{i}-K_{i})-\left(A_{L}-\sum_{i=1}^{L}K_{i}\right)$ (26)
$=\frac{1}{L}\sum_{i=1}^{L}A_{i}-A_{L}-\frac{1}{L}\sum_{i=1}^{L}K_{i}+\sum_{i=1}^{L}K_{i}$ (27)
$=\frac{1}{L}\sum_{i=1}^{L}A_{i}-A_{L}+\frac{L-1}{L}\sum_{i=1}^{L}K_{i}$ (28)
$=-\frac{1}{L}\sum_{i=1}^{L}(A_{L}-A_{i})+\frac{L-1}{L}\sum_{i=1}^{L}K_{i}$ (29)
Step 4: Apply the assumptions:
  • By Assumption 3: Each $A_{L}-A_{i}\geq 0$, so $\frac{1}{L}\sum_{i=1}^{L}(A_{L}-A_{i})\leq 0$
  • By Assumption 4: $\frac{L-1}{L}\sum_{i=1}^{L}K_{i}\leq\frac{1}{L}\sum_{i=1}^{L}(A_{L}-A_{i})$. Justification: On average across layers, the extra coding cost—quantified by the KL between $q$ and $p$, i.e., $\tfrac{L-1}{L}\sum_{i=1}^{L}K_{i}$—of shaping the latents must be no larger than the predictive gain—the increase in expected log-likelihood, i.e., $\tfrac{1}{L}\sum_{i=1}^{L}(A_{L}-A_{i})$. In short, depth’s benefit pays for its inference complexity.
Step 5: Conclude:
$\frac{1}{L}\sum_{i=1}^{L}\mathcal{E}_{i}-\mathcal{E}_{NN}=-\frac{1}{L}\sum_{i=1}^{L}(A_{L}-A_{i})+\frac{L-1}{L}\sum_{i=1}^{L}K_{i}\leq 0$ (30)
Therefore: $\frac{1}{L}\sum_{i=1}^{L}\mathcal{E}_{i}\leq\mathcal{E}_{NN}$
This completes the proof that the mean of layer-wise ELBOs provides a valid lower bound for the global ELBO.
A.3.2 Theorem 2: MI preservation of Random Projection
Let $h_{i}=f_{i}(h_{i-1})\in\mathbb{R}^{d}$ be the deterministic activation at layer $i$, and define the projected representation $v_{i}=R_{i}h_{i}$ using the fixed random matrix $R_{i}\sim\mathcal{N}(0,1/d^{\prime})$. Assume $h_{i}$ has bounded second moments. Let $n$ denote the effective number of distinct activation patterns in the support of the distribution of $h_{i}$." Then, for any $\varepsilon\in(0,1)$, if $d^{\prime}=O(\varepsilon^{-2}\log n)$, the mutual information satisfies, with high probability over $R_{i}$,
$I(x;v_{i})\geq I(x;h_{i})-\delta(\varepsilon),$ (31)
where $\delta(\varepsilon)\to 0$ as $\varepsilon\to 0$, and $x$ denotes the network input.
Proof. For any random variables $X$ and $Y$, the mutual information is defined as:
I(X;Y)=H(Y)H(Y|X)I(X;Y)=H(Y)-H(Y|X) (32)
Since $h_{i}$ is a deterministic function of $x$, we have $H(h_{i}|x)=0$, thus:
$I(x;h_{i})=H(h_{i})$ (33)
For the projected representation, given $x$, $v_{i}|x$ follows a Gaussian distribution with entropy:
$H(v_{i}|x)=\frac{d^{\prime}}{2}\log(2\pi e\frac{|h_{i}(x)|^{2}}{d^{\prime}})$ (34)
By the Johnson-Lindenstrauss lemma [19], for any two points $h_{i}(x_{1})$ and $h_{i}(x_{2})$, with high probability:
$(1-\varepsilon)\|h_{i}(x_{1})-h_{i}(x_{2})\|^{2}\leq\|v_{i}(x_{1})-v_{i}(x_{2})\|^{2}\leq(1+\varepsilon)\|h_{i}(x_{1})-h_{i}(x_{2})\|^{2}$ (35)
when $d^{\prime}=O(\varepsilon^{-2}\log n)$.
For Gaussian distributions with covariance $\Sigma=\frac{1}{d^{\prime}}I_{d^{\prime}}$, the KL divergence is:
$D_{KL}(p(v_{i}|x_{1})|p(v_{i}|x_{2}))=\frac{d^{\prime}}{2}\|R_{i}h_{i}(x_{1})-R_{i}h_{i}(x_{2})\|^{2}$ (36)
The preservation of distances directly implies preservation of distinguishability between different inputs. By discretizing the spaces of $h_{i}$ and $v_{i}$ and applying quantization theory, we can derive:
$I(x;v_{i})\geq(1-\varepsilon^{\prime})I(x;h_{i})-C\varepsilon^{\prime}\log(1/\varepsilon^{\prime})$ (37)
where $\varepsilon^{\prime}=O(\varepsilon)$ and $C$ is a constant depending on the bounded second moments of $h_{i}$.
Define $\delta(\varepsilon)=\varepsilon^{\prime}I(x;h_{i})+C\varepsilon^{\prime}\log(1/\varepsilon^{\prime})$. Since $I(x;h_{i})$ is bounded by our assumptions, $\delta(\varepsilon)\to 0$ as $\varepsilon\to 0$. Therefore, with high probability over the choice of $R_{i}$:
$I(x;v_{i})\geq I(x;h_{i})-\delta(\varepsilon)$ (38)
where $\delta(\varepsilon)\to 0$ as $\varepsilon\to 0$.
A.4 Model Architecture and Experiment Details
A.4.1 Computer Resources
All experiments were conducted on a single NVIDIA A100 GPU with 40GB of memory. No multi-GPU or distributed training was used.
A.4.2 Datasets
In this paper, we evaluate SVP on a range of benchmark datasets.
  • MNIST is a handwritten digit image dataset over 10 classes including 60,000 images for training and 10,000 images for testing. Each image is a 28 × 28 gray-scale image.
  • Fashion MNIST contains fashion items images such as clothing and shoes. It consists of a training set of 60,000 grayscale images and a test set of 10,000 images. Each image has a 28 × 28 size and is categorized into 10 classes.
  • CIFAR-10 consists of 32 × 32 RGB images for object recognition with 50,000 images for training and 10,000 images for testing. It has 10 classes, with 5,000 training and 1,000 testing images per class.
  • CIFAR-100 comprises a total of 60,000 32 × 32 RGB images distributed across 100 classes. Within each class, 500 images are allocated for training, while 100 images are for testing.
  • Tiny-ImageNet is a downsampled subset of ImageNet to a size of 64 × 64. This dataset consists of 200 classes and each class contains 500 images for training and 100 images for testing.
  • ImageNet-1Kis a large-scale image classification dataset containing over 1.28 million training images and 50,000 validation images across 1,000 classes. Images are typically resized to 224 × 224 pixels for training.
A.4.3 MLP
Architecture
For MNIST, we use a 2-layer MLP with 800 neurons per layer and ReLU activations, followed by dropout. For CIFAR-10 and CIFAR-100, we adopt a 3-layer MLP with 1000 neurons per layer, also using ReLU activations and dropout regularization.
Experimental Details.
During training, we apply random horizontal flipping and normalization as standard data augmentation. Models are optimized using the Adamax optimizer with a learning rate of 0.001, trained for 100 epochs across all datasets.
A.4.4 CNN
The architecture and training configurations used for CNN experiments are summarized in Table A5.
Dataset Network Architecture Training Details
FMNIST
SIZE: $\times28$
CLASS: 10
Conv64k3 - MaxPool2 - Conv128k3 - MaxPool2 - Conv256k3 - Conv256k3 - MaxPool2 - Conv512k3 - MaxPool2 - FC10 Data Aug: Normalize
Optimizer: Adam
Learning Rate: 0.003
Batch Size: 128
Epochs: 100
CIFAR10/100
SIZE: $\times32$
CLASS: 100
Conv256k3 - MaxPool2 - Conv512k3 - MaxPool2 - Conv512k3 - Conv1024k3 - MaxPool2 - Conv1024k3 - MaxPool2 - Conv1024k3 - Conv1024k3 - FC2048 - FC100 Data Aug: Crop, Flip, AutoAugment, Normalize
Optimizer: Adam
Learning Rate: 0.003
Batch Size: 128
Epochs: 500
Imagenette
SIZE: $\times224$
CLASS: 10
Conv64k3 - MaxPool2 - Conv128k3 - MaxPool2 - Conv256k3 - Conv256k3 - MaxPool2 - Conv512k3 - MaxPool2 - Conv512k3 - Conv1024k3 - FC10 Data Aug: Crop, Flip, Normalize
Optimizer: Adam
Learning Rate: 0.001
Batch Size: 128
Epochs: 500
Tiny-Imagenet
SIZE: $\times64$
CLASS: 200
Conv64k3 - MaxPool2 - Conv128k3 - MaxPool2 - Conv256k3 - Conv256k3 - MaxPool2 - Conv512k3 - MaxPool2 - Conv512k3 - Conv1024k3 - FC4096 - FC200 Data Aug: Crop, Flip, Normalize
Optimizer: Adam
Learning Rate: 0.001
Batch Size: 128
Epochs: 500
Table A5: Network architectures and training configurations for CNNs.
A.4.5 VIT
Details of the ViT architecture and experimental setup are provided in Table A6.
CIFAR10/100
SIZE: $\times32$
CLASS: 100
blocks=7,
heads=12,
mlp-ratio=2,
embedding=384
Data Aug: Crop, Flip, AutoAugment, Normalize
Optimizer: AdamW
Learning Rate: 0.003
Batch Size: 128
Epochs: 500
Imagenette
SIZE: $\times224$
CLASS: 10
blocks=7,
heads=12,
mlp-ratio=2,
embedding=384
Data Aug: Crop, Flip, Normalize
Optimizer: AdamW
Learning Rate: 0.001
Batch Size: 128
Epochs: 500
Imagenet
SIZE: $\times224$
CLASS: 1000
blocks=12,
heads=6,
mlp-ratio=3,
embedding=384
Data Aug: Crop, Flip, Normalize
Optimizer: AdamW
Learning Rate: 0.0005
Batch Size: 256
Epochs: 500
Table A6: Network architectures and training configurations for ViTs.